# Synthesis Framework for Agentic RL

### v1

@@ Abstract
Large language model (LLM)–based agents are increasingly deployed in complex, tool-augmented environments. While reinforcement learning (RL) provides a principled mechanism for such agents to improve through interaction, its effectiveness critically depends on the availability of structured training tasks. In many realistic settings, however, no such tasks exist—a challenge we term task scarcity, which has become a key bottleneck for scaling agentic RL. Existing approaches typically assume predefined task collections, an assumption that fails in novel environments where tool semantics and affordances are initially unknown. To address this limitation, we formalize the problem of Task Generation for Agentic RL, where an agent must learn within a given environment that lacks predefined tasks. We propose CuES, a Curiosity-driven and Environment-grounded Synthesis framework that autonomously generates diverse, executable, and meaningful tasks directly from the environment’s structure and affordances, without relying on handcrafted seeds or external corpora. CuES drives exploration through intrinsic curiosity, abstracts interaction patterns into reusable task schemas, and refines them through lightweight top-down guidance and memory-based quality control. Across three representative environments—AppWorld, BFCL, and WebShop—CuES produces task distributions that match or surpass manually curated datasets in both diversity and executability, yielding substantial downstream policy improvements. These results demonstrate that curiosity-driven, environment-grounded task generation provides a scalable foundation for agents that not only learn how to act, but also learn what to learn. The code is available at https://github.com/modelscope/AgentEvolver/research/CuES.
@@ 1 Introduction
Large language model (LLM)–based agents are increasingly deployed in complex, tool-augmented environments such as graphical user interfaces and open-ended web platforms (shang2025rstar2; li2025system). Reinforcement learning (RL) provides a principled mechanism for such agents to improve their policies through interaction and feedback (mai2025agent; lin2025comprehensive). However, in many realistic environments, no explicit training tasks are available. The agent can interact with the environment, but it lacks a structured set of tasks on which it can sample experience and perform RL optimization (cheng2024seeclick; zhang2024agentohana). As a result, this task scarcity has become a major bottleneck for scaling RL-based agent learning.
Existing approaches to agentic RL typically assume that a collection of training tasks is already provided (gao2024agentscope; yu2025demystifying; chinta2020agentic). These assumptions hold in controlled benchmarks but fail in novel environments, where tool semantics and affordances are initially unknown. In such cases, constructing a sufficiently diverse and executable set of tasks by hand is both costly and brittle, and the absence of such tasks leaves the agent without a foundation for effective learning (zhang2025agentrl; zeng2025glm). Consequently, the lack of training tasks—not the lack of algorithms—often limits progress in agentic RL(plaat2025agentic; singh2025agentic; sapkota2025ai).
To address this fundamental limitation (shi2025taskcraft; team2025kimi), we focus on the overlooked yet essential setting of training an agent within a given environment when no tasks are predefined. Specifically, we aim to answer the following question: Given an interactive environment but no predefined training tasks, how can an agent autonomously generate diverse, solvable, and useful tasks that enable effective policy learning? To this end, we first formulate and define the task generation problem for such environment-conditioned learning. We conceptualize task generation as the process of constructing a meaningful and solvable set of tasks directly from the environment’s structure and affordances, thereby bridging the gap between the environment and the agent’s learning process. This formulation establishes task generation as a first-class research problem in the development of self-improving LLM-based agents.
@@ 2 Related Work
To address this data scarcity, existing work generally follows two complementary lines. Top-down approaches start from human-written seed goals or LLM proposals and expand them into task instructions. This strategy benefits from clear goal specification and controllable semantics, but often decouples task generation from environment dynamics, producing instructions that look plausible in text yet fail during execution(li2024uieffects; lu2024weblinx; lai2024autowebglm). In contrast, bottom-up approaches explore the environment first and then abstract tasks from discovered trajectories, ensuring that proposed problems are feasible in situ. While this improves executability and behavioral grounding, it typically suffers from weak goal alignment, inefficient exploration, and domain-specific heuristics that limit transferability. Moreover, the current exploration of bottom-up approaches only targets a few narrow areas such as web search, and the implementation is also highly coupled to specific tasks, which makes it difficult to be applied in practice(sun2024osgenesis).
@@ 2.1 Top-down imitation-based synthesis: mimicking existing data
Top-down pipelines typically rely on human-authored tasks or expand a small seed goal set using LLMs(wu2025webdancer; li2025websailor; tao2025webshaper). Recent web and GUI agents illustrate both the promise and the pitfalls of this approach. Systems such as AutoWebGLM (lai2024autowebglm) and WebLINX (lu2024weblinx) operate with explicit high-level goals or dialogue instructions and often imitate patterns present in existing benchmarks. While this improves goal clarity, it tends to decouple proposals from environment dynamics, leading to low executability when intermediate preconditions are missing or tool affordances are misidentified. Empirical analyses show that scale alone does not resolve these issues: UI control studies find that simply increasing dataset size does not guarantee coherent, solvable trajectories (li2024uieffects), and self-improvement strategies without grounded verification may propagate subtle errors across steps (patel2024selfimprove).
The core limitations of imitation-based synthesis are twofold. First, dependence on predefined high-level tasks restricts scalability and curtails diversity, since the proposal space is ultimately bounded by seed goal patterns (lai2024autowebglm). Second, quality is hard to ensure: early-step mistakes or mismatched goals can corrupt entire trajectories, yielding incomplete or incoherent data even when text instructions look plausible (li2024uieffects; patel2024selfimprove). These issues form a bottleneck for advancing agents from scripted automation to robust autonomy in real environments.
@@ 2.2 Bottom-up exploration: discovering tasks from interaction
Recent research has attempted to synthesize data directly from environmental interactions under specific benchmarks, overturning the traditional process of ”specifying the task first, then collecting the trajectory”. OS-Genesis (sun2024osgenesis) exemplifies this reversal for GUI agents: an agent first explores the desktop and records raw traces, then retrospectively derives high-level tasks from the discovered behaviors, filtering trajectories to ensure quality and solvability. By grounding tasks in what the environment can actually support, this paradigm improves executability and increases variety relative to task-first scripting. These systems highlight a central advantage of bottom-up synthesis: proposals are feasible by construction because they originate from witnessed interactions.
Despite these gains, bottom-up pipelines face key challenges. Without sufficient guidance, exploration can drift, producing many trajectories that lack clear purpose or fail midway, which lowers overall yield and inflates redundancy (murty2024bagel). Moreover, many implementations are tailored to specific domains (e.g., GUI/web navigation), raising concerns about cross-domain generalization and transferability of the synthesized corpora (sun2024osgenesis). In practice, the absence of simple mechanisms to align discovery with desired coverage or stage-specific training needs can lead to substantial ineffective exploration and narrow gains outside the source domain.
@@ 5.2 Performance Comparison
We compare CuES with representative baselines (grattafiori2024llama; guo2024deepseek; yao2023react) on AppWorld, WebShop, and BFCL v3 Multi-Turn Base under the metrics of greedy.
@@
In Table 2, 3 and 4, we compare CUES against recent baselines on AppWorld, WebShop, and BFCL v3 Multi-Turn Base. We report avg@8 and greedy under the no-think and prompt configurations of Qwen2.5 and Qwen3. Since the baseline performance of the Qwen3 series on the BFCL v3 was lower than expected, we report the baseline results of Qwen3 under the thinking mode on the BFCL v3. On AppWorld, CUES attains 45.24% greedy, which closes to that of the closed-source GPT-4o. The gain is most pronounced on WebShop, where CUES reaches 64.10% greedy versus 50.96% for qwen3-235b-a22b. On BFCL v3 Multi-Turn Base, CUES achieves 44.15% greedy, edging out o3 at 44.0%. Looking at 1, CUES delivers a macro-average of roughly 51.2% greedy and 50.7% avg@8 across the three datasets, substantially higher than same-scale non-CUES backbones. These results indicate that environment-grounded synthesis with light top-down steering translates into consistent single-path and sampled execution gains across diverse domains. Notably, even when compared with models that have far more parameters or closed-source model than CUES (e.g., DeepSeek-V3(liu2024deepseek), GPT-5, o3(achiam2023gpt) and Gemini-2.5 Pro(comanici2025gemini)), CUES still outperforms by a large margin, underscoring the impact of our synthesis rather than raw scale.

### v2

@@ Abstract
Large language model (LLM)–based agents are increasingly deployed in complex, tool-augmented environments. While reinforcement learning (RL) provides a principled mechanism for such agents to improve through interaction, its effectiveness critically depends on the availability of structured training tasks. In many realistic settings, however, no such tasks exist—a challenge we term task scarcity, which has become a key bottleneck for scaling agentic RL. Existing approaches typically assume predefined task collections, an assumption that fails in novel environments where tool semantics and affordances are initially unknown. To address this limitation, we formalize the problem of Task Generation for Agentic RL, where an agent must learn within a given environment that lacks predefined tasks. We propose CuES, a Curiosity-driven and Environment-grounded Synthesis framework that autonomously generates diverse, executable, and meaningful tasks directly from the environment’s structure and affordances, without relying on handcrafted seeds or external corpora. CuES drives exploration through intrinsic curiosity, abstracts interaction patterns into reusable task schemas, and refines them through lightweight top-down guidance and memory-based quality control. Across three representative environments—AppWorld, BFCL, and WebShop—CuES produces task distributions that match or surpass manually curated datasets in both diversity and executability, yielding substantial downstream policy improvements. These results demonstrate that curiosity-driven, environment-grounded task generation provides a scalable foundation for agents that not only learn how to act, but also learn what to learn. The code is available at https://github.com/modelscope/AgentEvolver/tree/main/research/CuES.
@@ 1 Introduction
Large language model (LLM)–based agents are increasingly deployed in complex, tool-augmented environments such as graphical user interfaces and open-ended web platforms (Shang et al., 2025; Li et al., 2025b). Reinforcement learning (RL) provides a principled mechanism for such agents to improve their policies through interaction and feedback (Mai et al., 2025; Lin et al., 2025). However, in many realistic environments, no explicit training tasks are available. The agent can interact with the environment, but it lacks a structured set of tasks on which it can sample experience and perform RL optimization (Cheng et al., 2024; Zhang et al., 2024). As a result, this task scarcity has become a major bottleneck for scaling RL-based agent learning.
Existing approaches to agentic RL typically assume that a collection of training tasks is already provided (Gao et al., 2024; Yu et al., 2025; Chinta & Karaka, 2020). These assumptions hold in controlled benchmarks but fail in novel environments, where tool semantics and affordances are initially unknown. In such cases, constructing a sufficiently diverse and executable set of tasks by hand is both costly and brittle, and the absence of such tasks leaves the agent without a foundation for effective learning (Zhang et al., 2025; Zeng et al., 2025). Consequently, the lack of training tasks—not the lack of algorithms—often limits progress in agentic RL(Plaat et al., 2025; Singh et al., 2025; Sapkota et al., 2025).
To address this fundamental limitation (Shi et al., 2025; Team et al., 2025), we focus on the overlooked yet essential setting of training an agent within a given environment when no tasks are predefined. Specifically, we aim to answer the following question: Given an interactive environment but no predefined training tasks, how can an agent autonomously generate diverse, solvable, and useful tasks that enable effective policy learning? To this end, we first formulate and define the task generation problem for such environment-conditioned learning. We conceptualize task generation as the process of constructing a meaningful and solvable set of tasks directly from the environment’s structure and affordances, thereby bridging the gap between the environment and the agent’s learning process. This formulation establishes task generation as a first-class research problem in the development of self-improving LLM-based agents.
@@ 2 Related Work
To address this data scarcity, existing work generally follows two complementary lines. Top-down approaches start from human-written seed goals or LLM proposals and expand them into task instructions. This strategy benefits from clear goal specification and controllable semantics, but often decouples task generation from environment dynamics, producing instructions that look plausible in text yet fail during execution(Li et al., 2024; Lù et al., 2024; Lai et al., 2024). In contrast, bottom-up approaches explore the environment first and then abstract tasks from discovered trajectories, ensuring that proposed problems are feasible in situ. While this improves executability and behavioral grounding, it typically suffers from weak goal alignment, inefficient exploration, and domain-specific heuristics that limit transferability. Moreover, the current exploration of bottom-up approaches only targets a few narrow areas such as web search, and the implementation is also highly coupled to specific tasks, which makes it difficult to be applied in practice(Sun et al., 2024).
@@ 2.1 Top-down imitation-based synthesis: mimicking existing data
Top-down pipelines typically rely on human-authored tasks or expand a small seed goal set using LLMs(Wu et al., 2025; Li et al., 2025a; Tao et al., 2025). Recent web and GUI agents illustrate both the promise and the pitfalls of this approach. Systems such as AutoWebGLM (Lai et al., 2024) and WebLINX (Lù et al., 2024) operate with explicit high-level goals or dialogue instructions and often imitate patterns present in existing benchmarks. While this improves goal clarity, it tends to decouple proposals from environment dynamics, leading to low executability when intermediate preconditions are missing or tool affordances are misidentified. Empirical analyses show that scale alone does not resolve these issues: UI control studies find that simply increasing dataset size does not guarantee coherent, solvable trajectories (Li et al., 2024), and self-improvement strategies without grounded verification may propagate subtle errors across steps (Patel et al., 2024).
The core limitations of imitation-based synthesis are twofold. First, dependence on predefined high-level tasks restricts scalability and curtails diversity, since the proposal space is ultimately bounded by seed goal patterns (Lai et al., 2024). Second, quality is hard to ensure: early-step mistakes or mismatched goals can corrupt entire trajectories, yielding incomplete or incoherent data even when text instructions look plausible (Li et al., 2024; Patel et al., 2024). These issues form a bottleneck for advancing agents from scripted automation to robust autonomy in real environments.
@@ 2.2 Bottom-up exploration: discovering tasks from interaction
Recent research has attempted to synthesize data directly from environmental interactions under specific benchmarks, overturning the traditional process of ”specifying the task first, then collecting the trajectory”. OS-Genesis (Sun et al., 2024) exemplifies this reversal for GUI agents: an agent first explores the desktop and records raw traces, then retrospectively derives high-level tasks from the discovered behaviors, filtering trajectories to ensure quality and solvability. By grounding tasks in what the environment can actually support, this paradigm improves executability and increases variety relative to task-first scripting. These systems highlight a central advantage of bottom-up synthesis: proposals are feasible by construction because they originate from witnessed interactions.
Despite these gains, bottom-up pipelines face key challenges. Without sufficient guidance, exploration can drift, producing many trajectories that lack clear purpose or fail midway, which lowers overall yield and inflates redundancy (Murty et al., 2024). Moreover, many implementations are tailored to specific domains (e.g., GUI/web navigation), raising concerns about cross-domain generalization and transferability of the synthesized corpora (Sun et al., 2024). In practice, the absence of simple mechanisms to align discovery with desired coverage or stage-specific training needs can lead to substantial ineffective exploration and narrow gains outside the source domain.
@@ 5.2 Performance Comparison
We compare CuES with representative baselines (Grattafiori et al., 2024; Guo et al., 2024; Yao et al., 2023) on AppWorld, WebShop, and BFCL v3 Multi-Turn Base under the metrics of greedy.
@@
In Table 2, 3 and 4, we compare CUES against recent baselines on AppWorld, WebShop, and BFCL v3 Multi-Turn Base. We report avg@8 and greedy under the no-think and prompt configurations of Qwen2.5 and Qwen3. Since the baseline performance of the Qwen3 series on the BFCL v3 was lower than expected, we report the baseline results of Qwen3 under the thinking mode on the BFCL v3. On AppWorld, CUES attains 45.24% greedy, which closes to that of the closed-source GPT-4o. The gain is most pronounced on WebShop, where CUES reaches 64.10% greedy versus 50.96% for qwen3-235b-a22b. On BFCL v3 Multi-Turn Base, CUES achieves 44.15% greedy, edging out o3 at 44.0%. Looking at 1, CUES delivers a macro-average of roughly 51.2% greedy and 50.7% avg@8 across the three datasets, substantially higher than same-scale non-CUES backbones. These results indicate that environment-grounded synthesis with light top-down steering translates into consistent single-path and sampled execution gains across diverse domains. Notably, even when compared with models that have far more parameters or closed-source model than CUES (e.g., DeepSeek-V3(Liu et al., 2024), GPT-5, o3(Achiam et al., 2023) and Gemini-2.5 Pro(Comanici et al., 2025)), CUES still outperforms by a large margin, underscoring the impact of our synthesis rather than raw scale.